# Efficient stripe artefact removal by a variational method: application to light-sheet microscopy, FIB-SEM and remote sensing images

### v1

@@
Efficient stripe artefact removal by a variational method: application to light-sheet microscopy, FIB-SEM and remote sensing images
@@
journal: opticajournalarticletype: Research Article{abstract*}
Light-sheet fluorescence microscopy (LSFM) is used to capture volume images of biological specimens. It offers high contrast deep inside densely fluorescence labelled samples, fast acquisition speed and minimal harmful effects on the sample. However, the resulting images often show strong stripe artifacts originating from light-matter interactions. We propose a robust variational method suitable for removing stripes which outperforms existing methods and offers flexibility through two adjustable parameters. This tool is widely applicable to improve visual quality as well as facilitate downstream processing and analysis of images acquired on systems that do not provide hardware-based destriping methods. An evaluation of methods is performed on LSFM, FIB-SEM and remote sensing data, supplemented by synthetic LSFM images. The latter is obtained by simulating the imaging process on virtual samples.
@@ 1 Introduction
The observation by microscopy plays a fundamental role in understanding multicellular life and biological processes. Since the beginning of this century, light-sheet (fluorescence) microscopy (LSFM) has gained popularity in 3D-imaging of biological specimens. It provides high resolution and fast acquisition speeds [1] while minimizing harmful exposure to light [2]. However, the acquired images are often accompanied by strong stripe artifacts caused by light absorption and scattering, see Figure 1 for reference. The corruptions can occur in large parts of the imaged volume and obscure underlying structures, thus impairing their appearance and complicating further data analysis. There is a large quantity of articles concerned with stripe artifacts in LSFM [3, 4, 5, 6, 7, 8, 9]. In a recent paper, Ricci and colleagues [10] present a detailed summary of previous research on removing stripes from LSFM images. They discuss hardware solutions to prevent stripe formation, algorithms to remove stripes through post-processing and hybrid methods which combine ideas from both worlds. At first glance, the prevention of artifacts is clearly the preferential choice. However, this requires modifications of the imaging setup which are costly and often require entirely new microscopes. They may also limit the acquisition speed, the light-efficiency or be incompatible with certain advantageous image acquisition modes such as line-confocal acquisition [11, 12, 13]. In comparison, algorithmic post-processing is cheap and requires only computational power. Furthermore, corrupted images may already exist and specimens cannot be re-imaged, e.g., due to degradation and aging. For these reasons it is desirable to have powerful post-processing tools available for stripe removal.
@@
Figure 1: LSFM images corrupted by stripe artifacts. Cultivated cell organoid of mouse intestine cells (a), tail of Zebrafish larva (b) and cluster of HeLa cells (c).
Figure 2: Stripe artifacts in a FIB-SEM image slice of a tin bronze cast alloy (a) and Terra MODIS data (b).
@@
Stripe artifacts are not unique to LSFM. They are also common in various other imaging techniques such as atomic force microscopy (AFM) [14], focused ion beam scanning electron microscopy (FIB-SEM) [15] or remote sensing [16, 17], see Figure 2 for examples. Although having different causes, stripes are very similar in shape showing pronounced elongated structures in a common direction with mostly similar thickness of only few pixels. Their shape deviates strongly from underlying structures and presents a feature that can be used for detection and removal. In LSFM on the other hand, the appearance of stripes is much more diverse. While stripes are still aligned along a common direction, their length, thickness and intensity can vary drastically depending on the imaged structures. For example, a single stripe can be split into smaller separated segments of lengths equal to that of underlying structures, see Figure 1. Additionally, stripes display a much larger range of thickness and depth such that individual stripes can even appear in multiple slices of 3D volume images. This drastically complicates the correct detection and removal of such artifacts, especially since stripe removal methods heavily rely on the distinctive shape differences to image structures.
@@
In this paper, we discuss the two major categories of stripe removal methods in research, Fourier filtering and variational methods. We explain the functionality of Fourier filtering methods on the example of the multi-directional stripe remover (MDSR) [6] and present a modification to increase its performance. For the variational methods we propose a model which is based on the work of Liu and colleagues [16] and performs exceptionally well on all data tested. Furthermore, we suggest possible adaptations to make it applicable to a broad range of stripe removal settings. Afterwards, we evaluate and compare the performance of the methods using real and synthetic image data. The latter is obtained through physically correct simulation of light transport through randomly generated samples and provides a stripe-free reference image. The simulation is achieved using the Python package biobeam [18] and has to our knowledge not been used before to validate stripe removal. The availability of a ground truth which is unavailable in experimentally acquired images allows for a more precise assessment of the effectiveness of stripe removal methods as common quality metrics such as the peak signal-to-noise ratio (PSNR) or the multi-scale structural similarity index measure (MS-SSIM) can be applied. For the evaluation of real and synthetic data we also rely on visual inspection and the curtaining metric proposed by Roldán [19] which measures the amount of corruptions by stripes and does not require reference images.
@@ 2 Methods
The majority of methods proposed for stripe removal fall under one of two categories which we refer to as Fourier filtering and variational methods. In the following, we concentrate only on these categories, as they generalize well to several imaging methods and variations in the appearance of image structures and stripes. Additional approaches are average filtering [20, 21], histogram matching [22, 23], spline interpolation [24] and only recently neural networks [25, 26]. However, these are usually tailored for a specific appearance of images and stripes such that they are harder to transfer to other scenarios.
@@ 2.1 Fourier Filtering
The Fourier transform yields a mathematical decomposition of the image content into constituent frequencies. In image processing, the breakdown into a frequency spectrum is used to detect and modify periodic structures of certain frequencies or entire bands of frequencies. We assume that stripe artifacts point into the same direction, have small widths and appear repeatedly. Applying the Fourier transform, most information about stripes becomes encoded in frequency coefficients around a small band orthogonal to the stripe direction. On the other hand, natural structures tend to differ strongly from stripe artifacts and live on larger scales than the widths of stripes. Thus, image structure information is mostly dominated by low frequencies which are concentrated in central coefficients. This general behavior is visualized in Figure 3.
@@
By appropriate damping or removal of selected coefficients attributed to stripe artifacts, corruptions can be significantly reduced. However, these coefficients may also contain some information about structures. Therefore, it is essential to select coefficients carefully to minimally destroy structural information. This can be achieved, e.g., by using masked filters [27, 28] or a decision-based algorithm [14]. Nonetheless, using a prior decomposition by structure and scale through wavelets [29] or the non-subsampled contourlet transform (NSCT) [30] reduces the number of modified image structures more effectively. Since the general idea of Fourier filtering approaches remains the same across all methods, we only discuss the multi-directional stripe remover (MDSR) [6] based on the NSCT. However, we use a slightly modified filtering step to improve quality. The MDSR depicts an improvement over wavelet Fourier filtering (WFF) which was proposed by Münch et al. [29] and is commonly referenced and one of the earliest successful adoptions of Fourier filtering for removing stripes. The workflow of MDSR is visualized in Figure 4 and consists of the following steps:
@@
    Image Decomposition: Initially, stripe and image information are partially separated from each other by application of the NSCT. Using a pyramidal filter bank [30], the image is decomposed into subimages depending on direction and scale until a selected depth $n_{\mathrm{dec}}\in\mathbb{N}$ is reached. The number of decomposition directions is denoted by $n_{\mathrm{dir}}\in\mathbb{N}$.
@@
    Fourier Filtering: Subsequently, coefficients in the Fourier domain are modified and damped through filtering. This step is performed on subimages dependent on their decomposition directions $\theta_{i}\in[0,\pi],i=1,\dotsc,n_{\mathrm{dir}}$ such that directions close to stripe direction $\theta_{0}$ are strongly suppressed. Contrary to the proposition of Liang and colleagues [6], we apply damping only to subimages of directions with $|\theta_{i}-\theta_{0}|\leq\pi/4$. This reduces artifacting as shown in Figure 5. Considering vertical stripes, i.e. $\theta_{0}=\pi/2$, damping is performed via element-wise multiplication with
    $f_{i}(x,y)$ =$=\begin{cases}1-\exp\left(-\frac{(y-\frac{n_{y}}{2})^{2}}{2\sigma_{i}^{2}}\right)&\text{if }|\theta_{i}-\theta_{0}|\leq\pi/4,\\ 1&\text{otherwise},\end{cases}$
    $\sigma_{i}$ $=\sigma\exp\left(-0.5*\frac{(\theta_{0}-\theta_{i})^{2}}{\sigma_{a}^{2}}\right),$
    with Gaussian standard deviation $\sigma$ and vertical image size $n_{y}$. The $\sigma_{i}$ are adjustable damping parameters which depend on the deviation of directions $\theta_{i}$ to the stripe direction $\theta_{0}$. The parameter $\sigma_{a}>0$ describes how quickly damping is reduced when moving away from the stripe direction.
@@
    Backtransformation: Finally, the modified Fourier coefficients are transformed back into individual subimages. These are then recombined using NSCT reconstruction to yield a destriping result.
@@
Fourier filtering approaches such as MDSR rely on the condensation of stripe information in the Fourier domain and separability from image information. These prerequisites are given when stripes and structures have different shapes and appearances. However, stripes in LSFM images cover a wide range of lengths, widths and intensities. This results in stripe information to be less concentrated around the horizontal coefficient band and more mixed up with image information. Therefore, stripe removal becomes more difficult and can lead to the introduction of image artifacts through modification of image information.
@@ 2.2 Variational Methods
Convex functions possess special properties that make them ideal for optimization problems. Under mild assertions, convex functions have unique minimizers which can be approximated using well-studied optimization algorithms, see [31]. Furthermore, convexity is an additive property of functions that allows for capturing multiple aspects of the stripe removal problem in a single objective function. Building on this concept, variational methods use a problem-specific weighted function consisting of convex building blocks, most commonly norms. The individual blocks penalize unwanted features of the clean image $u$ and stripes $s$ through positive contribution to the function. A corresponding minimizer should then reasonably fulfill the desired properties, i.e., optimizing the objective function yields a solution to the problem. This approach is well known in image processing and commonly used for denoising, segmentation and active contours, see, e.g, [32]. To give further insight into the construction of a suitable objective function, we motivate our model in the following. It has been originally developed by adjusting the functional proposed by Fitschen et al. [15] for curtaining noise in FIB-SEM images to the LSFM setting. The result shows resemblance to the proposition by Liu et al. [16] but extends applicability to 3D images and includes an additional regularization term which improves performance and consistency.
@@ 2.2.1 Construction
Consider the corrupted image $u_{0}\in[0,1]^{N}$, the clean image $u\in[0,1]^{N}$ and the stripe image $s\in[-1,1]^{N}$ where $N=n_{x}\times n_{y}\times n_{z}$. We assume that $u_{0}=u+s$ holds true, i.e., the corrupted image can be expressed as sum of a clean image and stripes. Let $\nabla_{x},\nabla_{y}$ and $\nabla_{z}$ denote the directional difference operators for the three coordinate directions and assume that stripes point in $y$-direction.
The objective function is constructed as superposition of individual convex building blocks which encode certain properties of the clean image and stripes. These are generally grouped into a data term $D(u)$ and a noise term $N(s)$ such that the general optimization problem reads
@@
is a frequently used data term in image processing, see [33, 16]. It expresses that natural images are piece-wise smooth, i.e., they contain only few strong edges, which is reflected by overall small directional differences in all directions. In contrast, stripe artifacts are usually thin and elongated. This corresponds to large values of the difference operators $\nabla_{x}s$ and $\nabla_{z}s$ orthogonal to the stripe direction and small values of the operator $\nabla_{y}s$ in direction parallel to it. This behaviour is visualized in Figure 6 and incorporated by the function
@@
Combined with the total variation (2), these capture the essence of stripe removal in a single data and noise term. However, for increased performance regularization is added through
@@
Function (4) is $\ell_{1}$-regularization and promotes sparsity in the stripe image. It reflects the generally small area affected by artifacts and avoids unnecessary detection of image structures as stripes. On the contrary, (5) is included to guarantee the outcome to live in the same value range as the input. Specifically, values remain non-negative which reflects the natural interpretation of the signal. Hence, further post-processing such as adjusting contrast or the intensity profile becomes obsolete. With (2)-(5) the resulting optimization problem reads
@@ 2.2.2 Minimization
Convex optimization problems such as (6) can be solved via the primal-dual gradient hybrid method with extrapolation of the dual variable (PDHGMp), see [31]. In principle, the iterative algorithm relies on alternating between minimizing the primal and maximizing the dual problem. Under mild assertions on the objective function, the sequence generated by PDHGMp converges to a solution. More details on the PDHGMp and alternatives can be found in [31, 34].
The weighting parameters $\mu_{1},\mu_{2}$ in (6) provide adjustability to control different aspects of the stripe removal, enabling the production of a desired result, see Appendix B. The following rules of thumb can be established:
@@
    $\mu_{1}$ controls the desired smoothness of the outcome. It strongly regulates the amount of stripe removal to be applied. Larger values will result in stronger removal of stripe artifacts. However, vertical and stripe-like image structures may become affected which results in a smoothed image.
@@
    $\mu_{2}$ acts like a counterpart to $\mu_{1}$ and controls how strictly attention is paid to the stripe-like appearance of affected structures. Increasing its value generally tends towards a lower reduction of artifacts and image structures. However, in combination with $\mu_{1}$ it enforces that for the most part only structures with concise stripe properties, e.g., thin, long and vertically aligned, are removed. Structures which deviate from this become less likely to be affected.
@@
    The ratio of $\mu_{1}$ and $\mu_{2}$ is crucial and has a strong influence on the general outcome. Scaling both parameters by an equal factor changes the amount of stripe removal while keeping the effect on image structures mostly unchanged.
In addition to offering adjustability through two parameters, variational methods are generally less susceptible to stripes deviating from being thin and long which we will see in Section 3. Furthermore, they can easily be adapted towards specific settings. For example, considering 2D image data we replace the total variation (2) by its lower-dimensional counterpart
@@
Alternatively, it is common for LSFM images to have a lower spatial resolution along the displacement direction $z$ of the light-sheet which can be incorporated by using $\nabla^{*}_{z}u=\rho_{z}\nabla_{z}u$ with $0\leq\rho_{z}\leq 1$. This emphasizes that information along the $z$-direction is less reliable and coherent. The case $\rho_{z}=1$ corresponds to (6) and $\rho_{z}=0$ to the 2D case. We can also consider oblique stripe artifacts as described in [16] by replacing $\nabla_{y}s$ with a suitable directional difference operator $\nabla_{\theta}s$ in (3). Lastly, a simultaneous multi-directional stripe removal is possible using a sum of penalization terms $\left\lVert\nabla_{\theta_{i}}s\right\rVert_{1}$ of different stripe directions $\theta_{i}$.
2.2.3 State of the art
Besides our proposition, there exist a variety of different objective functions which are usually tailored for specific cases. An early and more general proposition comes from Fehrenbach and colleagues [35] designed to remove repeating artifacts. The model relies on the representation of artifacts as
$s=\sum_{i=1}^{m}\lambda_{i}\ast\psi_{i},\quad m\in\mathbb{N}$
with elementary noise patterns $\psi_{i}$ and $\lambda_{i}$ containing position information. This approach makes it applicable to not only stripes but any kind of noise created by repeating specific patterns. In a follow-up paper by Escande et al. [36], multiplicative noise of the same form is considered. This represents stripe formation in LSFM more closely. However, the authors have noticed a significant change in contrast as side effect of their method which can strongly alter the image appearance. Furthermore, the selection of reasonable elementary patterns $\psi_{i}$ is not straightforward since large variations in stripe sizes are observed in LSFM. Other propositions include [17] utilizing a low-rank assumption of $s$ and [15] for the removal of curtaining artifacts including stripes. While these are prime examples of variational methods, their assumptions deviate strongly from observations in LSFM.
@@ 2.2.3 State of the art
Another proposition for stripe removal stems from Liu and colleagues [16] who proposed a functional for oblique stripe artifacts which corresponds to (6) for vertical stripes but without the indicator function $\iota_{[0,1]^{N}}(u)$ to restrict values to the range $[0,1]$. Additionally, they only consider the 2D-total variation of (2). In Figure 7 we display stripe removal results to visually compare the influence of processing in 2D and 3D and the difference of restricting optimization to the value range $[0,1]$ or not. At first glance, results are almost identical in terms of their visual appearance. However, at closer inspection some noticeable differences can be spotted. For example, when moving from 2D to 3D, we see a subtle increase in brightness in the background area left to the center. Inspection of the 3D image stack reveals that structural components can be found in the neighbouring slices which explains this phenomenon. Therefore, when interested in a single image slice it is sufficient to use the 2D variation. However, for 3D the incorporation of depth information ensures reasonably smooth intensity profiles and prevents sudden jumps. This becomes beneficial when further processing, such as image segmentation, is performed on the 3D image. On the other hand, the indicator function does not cause drastically different results if we clip the intensity profile to the same range after processing. However, as the affected areas are not negligible, we argue that removing the need for further steps is beneficial and prevents the introduction of errors.

### v2

# Effective stripe artifact removal by a variational method: application to light-sheet microscopy, FIB-SEM and remote sensing images

@@
Effective stripe artifact removal by a variational method: application to light-sheet microscopy, FIB-SEM and remote sensing images
@@
journal: opticajournalarticletype: Research Article
{abstract*}
Light-sheet fluorescence microscopy (LSFM) is used to capture volume images of biological specimens. It offers high contrast deep inside densely fluorescence labelled samples, fast acquisition speed and minimal harmful effects on the sample. However, LSFM images often show strong stripe artifacts originating from light-matter interactions. We propose a robust variational method suitable for removing stripes which outperforms existing methods and offers flexibility through two adjustable parameters. This tool is widely applicable to improve visual quality as well as facilitate downstream processing and analysis of images acquired on systems that do not provide hardware-based destriping methods. An evaluation of methods is performed on LSFM, focused ion beam scanning electron microscopy (FIB-SEM) and remote sensing data, supplemented by synthetic LSFM images. The latter is obtained by simulating the imaging process on virtual samples.
@@ 1 Introduction
The observation by microscopy plays a fundamental role in understanding multicellular life and biological processes. Since the beginning of this century, light-sheet (fluorescence) microscopy (LSFM) has gained popularity in 3D-imaging of biological specimens. It provides high resolution and fast acquisition speeds [1] while minimizing harmful exposure to light [2]. However, the acquired images are often accompanied by strong stripe artifacts caused by light absorption and scattering, see Fig. 1 (a-c). The corruptions can occur in large parts of the imaged volume and obscure underlying structures, thus impairing their appearance and complicating further data analysis. There is a large quantity of articles concerned with stripe artifacts in LSFM [3, 4, 5, 6, 7, 8, 9]. In a recent paper, Ricci and colleagues [10] present a detailed summary of previous research on removing stripes from LSFM images. They discuss hardware solutions to prevent stripe formation, algorithms to remove stripes through post-processing and hybrid methods which combine ideas from both worlds.
At first glance, the prevention of artifacts is clearly the preferential choice. However, this requires modifications of the imaging setup which are costly and often require entirely new microscopes. They may also limit the acquisition speed, the light-efficiency or be incompatible with advantageous image acquisition modes such as line-confocal acquisition [11, 12, 13]. In comparison, algorithmic post-processing is cheap and requires only computational power. Furthermore, corrupted images may already exist and specimens cannot be re-imaged, e.g., due to degradation and aging. For these reasons it is desirable to have powerful post-processing tools for stripe removal available.
@@
Figure 1: Images corrupted by stripe artifacts from LSFM (a-c), FIB-SEM (d) and remote sensing (e). Cultivated cell organoid of mouse intestine cells (a), tail of Zebrafish larva (b), cluster of HeLa cells (c), image slice of a tin bronze cast alloy (d) and Terrra MODIS data (e).
@@
Stripe artifacts are not unique to LSFM. They are also common in imaging techniques such as atomic force microscopy (AFM) [14], focused ion beam scanning electron microscopy (FIB-SEM) [15] or remote sensing [16, 17], see Fig. 1 (d-e) for examples. Although having different causes, stripes are very similar in shape showing pronounced elongated structures in a common direction with mostly similar thickness of only a few pixels. Their shape deviates strongly from underlying structures and presents a feature that can be used for detection and removal. In contrast, the appearance of stripes in LSFM is much more diverse. While stripes are still aligned along a common direction, their length, thickness and intensity can vary drastically depending on the imaged structures. For example, a single stripe can be split into smaller separated segments of lengths equal to that of underlying structures, see Fig. 1 (a). Additionally, stripes display a much larger range of thickness and depth such that individual stripes can appear in multiple slices of 3D volume images. This drastically complicates the correct detection and removal of such artifacts, as stripe removal methods heavily rely on the distinctive shape differences to image structures.
@@
In this paper, we discuss the two major categories of stripe removal methods in research, Fourier filtering and variational methods. We explain the functionality of Fourier filtering methods on the example of the multi-directional stripe remover (MDSR) [6] and present a modification to increase its performance. For the variational methods we propose a model which is based on the work of Liu and colleagues [16] and performs exceptionally well on all data tested. Furthermore, we suggest possible adaptations to make it applicable to a broad range of stripe removal settings. We evaluate and compare the performance of the methods using real and synthetic image data. The latter is obtained through physically correct simulation of light transport through randomly generated samples and provides a stripe-free reference image. The simulation is achieved using the Python package biobeam [18] and, to our knowledge, has not previously been used to validate stripe removal. The availability of a ground truth, which is unavailable in experimentally acquired images, allows for an objective validation of stripe removal methods as quality metrics such as the peak signal-to-noise ratio (PSNR) or the multi-scale structural similarity index measure (MS-SSIM) can be applied. For the evaluation of real and synthetic data we rely on visual inspection and the curtaining metric proposed by Roldán [19] which measures the amount of corruptions by stripes and does not require reference images.
@@ 2 Methods
The majority of methods proposed for stripe removal fall under the two categories Fourier filtering and variational methods. We will concentrate on these categories, as they generalize well to several imaging methods and variations in image structures and stripe appearance. Additional approaches are average filtering [20, 21], histogram matching [22, 23], spline interpolation [24] and recently neural networks [25, 26, 27, 28]. However, these are usually tailored for a specific appearance of images and stripes such that they are harder to transfer to other scenarios.
@@ 2.1 Fourier Filtering
The Fourier transform yields a mathematical decomposition of the image content into constituent frequencies. In image processing, the breakdown into a frequency spectrum is used to detect and modify periodic structures of certain frequencies or entire bands of frequencies. We assume that stripe artifacts point into the same direction, have small widths and appear repeatedly. In the Fourier domain, stripe information is encoded in frequency coefficients around a small band orthogonal to the stripe direction. In contrast, natural structures live on larger scales than the widths of stripes. Thus, image structure information is dominated by low frequencies which are concentrated in central coefficients. This general behavior is visualized in Fig. 2.
@@
By appropriate damping or removal of coefficients attributed to stripe artifacts, corruptions can be significantly reduced. This can be achieved, e.g., by using masked filters [29, 30, 31] or a decision-based algorithm [14]. A prior decomposition by structure and scale through wavelets [32] or the non-subsampled contourlet transform (NSCT) [33] reduces the effect of stripe removal on the image structure. We will use the multi-directional stripe remover (MDSR) [6] based on the NSCT. However, we use a slightly modified filtering step to improve quality. The MDSR depicts an improvement over wavelet Fourier filtering proposed by Münch et al. [32] which is commonly referenced and one of the earliest successful adoptions of Fourier filtering for removing stripes. The workflow of MDSR is visualized in Fig. 3 and consists of the following steps:
@@
    Image Decomposition: Initially, stripe and image information are partially separated by application of the NSCT. Using a pyramidal filter bank [33], the image is decomposed into subimages depending on direction and scale until a selected depth $n_{\mathrm{dec}}\in\mathbb{N}$ is reached. The number of decomposition directions is denoted by $n_{\mathrm{dir}}\in\mathbb{N}$.
@@
    Fourier Filtering: Subsequently, coefficients are filtered on subimages in the Fourier domain. Contrary to the original MDSR [6], we apply damping only to subimages of directions with $|\theta_{i}-\theta_{0}|\leq\pi/4$. This reduces artifacts as shown in Fig. 4. For vertical stripes, i.e., $\theta_{0}=\pi/2$, damping is performed via element-wise multiplication with
    $f_{i}(x,y)=\begin{cases}1-\exp\left(-\frac{(y-\frac{n_{y}}{2})^{2}}{2\sigma_{i}^{2}}\right)&\text{if }|\theta_{i}-\theta_{0}|\leq\pi/4,\\ 1&\text{otherwise},\end{cases},\quad\sigma_{i}=\sigma\cdot\exp\left(-\frac{(\theta_{0}-\theta_{i})^{2}}{2\sigma_{a}^{2}}\right),$
    with Gaussian standard deviation $\sigma$ and vertical image size $n_{y}$. The damping parameters $\sigma_{i},\,i=1,\ldots,n_{\mathrm{dir}},$ depend on the deviation of $\theta_{i}$ from the stripe direction $\theta_{0}$. The parameter $\sigma_{a}>0$ describes how quickly damping is reduced when moving away from the stripe direction.
@@
    Backtransformation: The modified Fourier coefficients are transformed back into individual subimages. NSCT reconstruction then yields the destriping result.
@@
Fourier filtering approaches such as MDSR rely on separation of stripe information and image content in the Fourier domain. However, stripes in LSFM images cover a wide range of lengths, widths and intensities such that stripe information is less concentrated around the horizontal coefficient band and more mixed up with image information. This complicates stripe removal.
@@ 2.2 Variational Methods
Variational methods achieve destriping by minimizing a problem-specific function consisting of a weighted sum of convex functions capturing multiple aspects of the stripe removal problem. Under mild assertions, convex functions have unique minimizers which can be approximated using well-studied optimization algorithms, see [34]. The individual functions penalize unwanted features of the clean image $u$ and stripes $s$ through positive contribution to the function. A corresponding minimizer should then reasonably fulfill the desired properties, i.e., optimizing the objective function yields a solution to the problem. This approach is well known in image processing and commonly used for denoising, segmentation and active contours [35]. We adjusted a functional originally proposed by Fitschen et al. [15] for curtaining noise in FIB-SEM images to the LSFM setting. The result shows resemblance to the proposition by Liu et al. [16] but is also applicable to 3D images and includes an additional regularization term which improves performance and consistency.
Besides our proposition, there exist a variety of alternative objective functions. An early and more general proposition comes from Fehrenbach and colleagues [36] designed to remove repeating artifacts. The model relies on the representation of artifacts in terms of elementary noise patterns which makes it applicable to any kind of noise created by repeating specific patterns. In a follow-up paper by Escande et al. [37], multiplicative noise of the same form is considered. This represents stripe formation in LSFM more closely. However, the authors have noticed a significant change in contrast as side effect of their method which can strongly alter the image appearance. Other propositions include [17, 38, 39, 40] utilizing sparsity and low-rank assumptions on $s$ and [41] with a combination of Fourier filtering and variational methods.
@@ 2.2.1 Construction
Consider the corrupted image $u_{0}\in[0,1]^{N}$, the clean image $u\in[0,1]^{N}$ and the stripe image $s\in[-1,1]^{N}$ where $N=n_{x}\times n_{y}\times n_{z}$. We assume that $u_{0}=u+s$. Let $\nabla_{x},\nabla_{y}$ and $\nabla_{z}$ denote the directional difference operators for the three coordinate directions and assume that stripes point in $y$-direction.
The objective function is grouped into a data term $D(u)$ and a noise term $N(s)$ such that the general optimization problem reads
@@
is a frequently used data term in image processing, see [42, 16]. The clean image $u$ is supposed to be piece-wise smooth. That is, it contains only few strong edges, which is reflected by overall small directional differences in all directions. In contrast, stripe artifacts are characterized by large values of the difference operators $\nabla_{x}s$ and $\nabla_{z}s$ orthogonal to the stripe direction and small values of the operator $\nabla_{y}s$ in direction parallel to it. This behaviour is visualized in Fig. 5 and incorporated by the function
@@
For increased performance regularization is added through
@@
$\ell_{1}$-regularization through (4) promotes sparsity in the stripe image. It reflects the generally small area affected by artifacts. The indicator (5) is included to guarantee the outcome to live in the same value range as the input. In particular, values remain non-negative. Hence, further post-processing such as adjusting contrast or the intensity profile becomes obsolete. The resulting optimization problem reads
@@ 2.2.2 Minimization
Convex optimization problems such as (6) can be solved via the primal-dual gradient hybrid method with extrapolation of the dual variable (PDHGMp), see [34]. Under mild assertions on the objective function, the sequence generated by PDHGMp converges to a solution, see [34, 43] for details.
The weighting parameters $\mu_{1},\mu_{2}$ in (6) provide adjustability to control different aspects of the stripe removal, see Supplement 1. The following rules of thumb can be established:
@@
    $\mu_{1}$ controls the smoothness of the outcome. It strongly regulates the amount of stripe removal to be applied. Larger values will result in stronger removal of stripe artifacts. However, vertical and stripe-like image structures may become affected which results in a smoothed image.
@@
    $\mu_{2}$ acts like a counterpart to $\mu_{1}$ and controls how strictly attention is paid to the stripe-like appearance of affected structures. Increasing its value generally tends towards a lower reduction of artifacts and image structures. However, in combination with $\mu_{1}$ it enforces that for the most part only structures with concise stripe properties, e.g., thin, long and vertically aligned, are removed.
@@
    The ratio of $\mu_{1}$ and $\mu_{2}$ has a strong influence on the general outcome. Scaling both parameters by an equal factor changes the amount of stripe removal while keeping the effect on image structures mostly unchanged.
In addition to offering adjustability through two parameters, variational methods are generally less susceptible to stripes deviating from being thin and long which we will see in Section 3. Furthermore, they can easily be adapted to specific settings. For example, for 2D image data we replace the total variation (2) by its lower-dimensional counterpart
@@
The lower spatial resolution of LSFM images along the displacement direction $z$ of the light-sheet can be incorporated by using $\nabla^{*}_{z}u=\rho_{z}\nabla_{z}u$ with $0\leq\rho_{z}\leq 1$. This emphasizes that information along the $z$-direction is less reliable and coherent. The case $\rho_{z}=1$ corresponds to (6) and $\rho_{z}=0$ to the 2D case. For oblique stripe artifacts as described in [16] we replace $\nabla_{y}s$ with a suitable directional difference operator $\nabla_{\theta}s$ in (3). Lastly, a simultaneous multi-directional stripe removal is possible using a sum of penalization terms $\left\lVert\nabla_{\theta_{i}}s\right\rVert_{1}$ of different stripe directions $\theta_{i}$.
@@
In Fig. 6 we visually compare the influence of processing in 2D and 3D and the difference of restricting optimization to the value range $[0,1]$ or not. At first glance, results are almost identical. However, at closer inspection some noticeable differences can be spotted. For example, when moving from 2D to 3D, we see a subtle increase in brightness in the background area left to the center. Inspection of the 3D image stack reveals that structural components can be found in the neighbouring slices which explains this phenomenon. Therefore, when interested in a single image slice it is sufficient to use the 2D variation. However, for 3D the incorporation of depth information ensures reasonably smooth intensity profiles and prevents sudden jumps. This becomes beneficial when further processing, such as image segmentation, is performed on the 3D image.
The indicator function does not cause drastically different results if we clip the intensity profile to the same range after processing. However, as the affected areas are not negligible, we argue that removing the need for further steps is beneficial and prevents the introduction of errors.